Diane is 23 years older than her daughter Amy. In six years, Diane will be twice as old as Amy. How old are they now.
Answer by Timnewman(323)

Let her daughter' age be "x"
Then Diane =x+23.
In six years;
Her daughter =x+6
Diane's age =(x+29)
But Diane at six years will be twice as her daughter Amy
Therfore,
x+29=2(x+6)
x+29=2x+12
2x-x=29-12
x=17
Amy's age = 17yrs
Daine's age =x+23
=17+23
=40yrs
HOPE THIS HELPS,
Best regards,
Timnewman::